OpenStudy (anonymous):

Part 1 [2 points] Is the following equation true? Part 2 [6 points] Use complete sentences to explain the properties used in making your decision. 3log2 x + .5log2 y - 3log 2z = log2x-cubed times square root of y all over z-cubed.

OpenStudy (dumbcow):

\[3\log x + .5\log y -3\log z \] \[\log x^{3} + \log \sqrt{y} - \log z^{3}\] \[\log (\frac{x^{3}\sqrt{y}}{z^{3}})\]